OpenStudy (twizttiez):

The spherical container has a radius of 21 cm. What is the approximate volume of the container? Use 3.14 to approximate pi and express your answer in hundredths. 38,772.72 cm^3

OpenStudy (twizttiez):

@sleepyjess sorry OS is messing up

OpenStudy (twizttiez):

But I think that's my answer

OpenStudy (acxbox22):

V=(4/3)(pi)(r^3) V=(4/3)(pi)(21^3) plug that into a calculator and tell me what u get your answer is close but thats not what i got

OpenStudy (twizttiez):

38772.72

OpenStudy (twizttiez):

@acxbox22

OpenStudy (anonymous):

Your answer is correct! Exactly 38772.72 cm^3

OpenStudy (twizttiez):

Thanks